OpenStudy (anonymous):

what is the slope if 8-2/2-2?

OpenStudy (anonymous):

Vertical

OpenStudy (anonymous):

Use parenthesis! 8-2/2-2 = 8-1-2 = 5 (8-2)/(2-2) = 6/0 = Undefined Very different slopes.